OpenStudy (anonymous):

Which of the following is equivalent to log8 16? Answer Choices: 2 1.248 4/3 3/4

OpenStudy (anonymous):



OpenStudy (raden):

solve for x 8^x = 16

OpenStudy (anonymous):

\[\Large \log_{8} 16\] answers the question \(\Large 8^? = 16\)

OpenStudy (anonymous):

Check if that is true for each of your options. \(\Large 8^2 = 16\) True or false?

OpenStudy (anonymous):

So would it end up being 2?

OpenStudy (anonymous):

You tell me. Is \(\Large 8^2 =16\)?

OpenStudy (raden):

8^x = 16 (2^3)^x = 2^4 2^(3x) = 2^4 3x = 4 nah, this should make easier for u now

OpenStudy (anonymous):

Thankss!

OpenStudy (raden):

welcome